Slope-intercept form of x + 4y - 4 = 0
Answer by umail08(87)

Slope-intercept form of x + 4y - 4 = 0
Y=mx+b …. Slope-intercept form .. m=slope and b =intercept
x + 4y – 4 = 0
4y = -x + 4 ….divide all terms by 4
y = -1/4x + 1 …. Y =mx + b